# Grey-body factors of higher dimensional regular black holes in quasi-topological theories

### v1

@@ I Introduction
Black holes are not perfect black bodies. Quantum mechanically, they emit radiation with a thermal spectrum modulated by frequency-dependent transmission coefficients, known as grey-body factors, which encode the interaction of the emitted quanta with the curved spacetime geometry outside the event horizon. These factors determine both the energy emission rates and the detailed spectral properties of Hawking radiation Hawking (1975), thereby providing a direct link between semiclassical quantum effects and the classical scattering problem for perturbations in black hole backgrounds. As such, grey-body factors play a central role in understanding black hole evaporation, stability, and potential observational signatures of modifications to General Relativity Page (1976a, b); Kanti (2004).
@@
In recent years, regular black holes – spacetimes free of curvature singularities – have attracted considerable attention as viable effective descriptions of quantum-corrected gravitational collapse. In many constructions, regularity is achieved either through effective matter sources or via higher-curvature corrections, while preserving the existence of an event horizon. From a phenomenological perspective, such models provide a setting in which quantum-gravity-inspired modifications can be explored without encountering pathological divergences at the classical level. However, regularity alone does not guarantee that the dynamical, thermodynamical, or radiative properties of these black holes closely resemble those of their singular counterparts. This makes a detailed analysis of wave propagation, scattering, and Hawking emission indispensable. Consequently, a substantial body of literature has been devoted to the study of perturbations, quasinormal spectra, grey-body factors, and Hawking radiation of various regular black hole models Simpson (2021); López and Ramírez (2023); Dutta Roy and Kar (2022); Meng and Zhang (2023); Li (2023); Konoplya et al. (2023a); Konoplya (2023a); Myrzakulov et al. (2023); Filho (2024); Bolokhov et al. (2025); Al-Badawi (2023); Konoplya et al. (2023b); Konoplya (2025); Jha (2023); Balart et al. (2023); Huang et al. (2024); Bolokhov (2024a); Al-Badawi and Kraishan (2024); Guo et al. (2024); Zhang et al. (2024); Pedrotti and Vagnozzi (2024); Gingrich (2024); Bolokhov and Konoplya (2025); Skvortsova (2025a); Bonanno et al. (2025); Skvortsova (2025b); Konoplya and Zhidenko (2025a); Bolokhov (2025); Bolokhov and Skvortsova (2025a, b); Lütfüoğlu et al. (2025a).
@@
Higher-curvature extensions of General Relativity are particularly well suited for such investigations. Among them, quasi-topological gravity occupies a distinguished position, as it allows for nontrivial higher-order curvature corrections while retaining second-order field equations for static, spherically symmetric spacetimes in arbitrary dimensions Boulware and Deser (1985); Wiltshire (1988); Cai and Guo (2004). When combined with regular black hole metrics, quasi-topological gravity provides a consistent and tractable framework in which higher-curvature effects and regular cores jointly influence black hole radiation and dynamical properties Bueno et al. (2025a); Konoplya and Zhidenko (2024a, b); Bueno et al. (2025b, c); Frolov and Zelnikov (2026). Notice, that black holes in theories with usual Gauss-Bonnet or Lovelock corrections are usually singular and their quasinormal spectrum has been extensively studied (see, for instance Abdalla et al. (2005); Moura and Schiappa (2007); Daghigh et al. (2007); Zhidenko (2008); Yoshida and Soda (2016); Cuyubamba et al. (2016); Konoplya and Zhidenko (2017a, b); Kokkotas et al. (2017); Konoplya and Zhidenko (2017c); Gonzalez et al. (2018); Zinhailo (2019); Carson and Yagi (2020); Konoplya and Zinhailo (2020a); Cano et al. (2020); Blázquez-Salcedo et al. (2020); Konoplya et al. (2020); Blázquez-Salcedo et al. (2022); Pierini and Gualtieri (2022); Konoplya (2023b) and references therein).
@@
In this work, we investigate grey-body factors and energy emission rates for higher-dimensional regular black holes arising in quasi-topological gravity. By solving the wave equations for the Maxwell field propagating in these backgrounds, we analyze how spacetime dimensionality and regularization parameter affect transmission probabilities and Hawking spectra. Special attention is devoted to deviations from purely thermal emission and to the role of the effective potential barrier surrounding the black hole. While Hawking radiation for two specific regular black hole solutions in quasi-topological gravity was previously considered in Ref. Konoplya and Zhidenko (2025b) for particular values of the coupling parameter, a comprehensive analysis covering a broader class of regular solutions and the full range of parameters has not yet been performed. The present work aims to fill this gap through a systematic study of grey-body factors and energy emission rates for regular black holes in quasi-topological gravity.
@@
In addition, we explore the correspondence between quasinormal modes and grey-body factors, treating it as a complementary aspect rather than a primary assumption. Quasinormal modes govern the characteristic damped oscillations of black holes and are closely related to the poles of scattering amplitudes in the complex frequency plane. It has been suggested in various contexts that features of the grey-body spectrum, such as characteristic frequency scales or transmission resonances, may be linked to the quasinormal mode spectrum Malik (2025a); Han and Gwak (2025); Lütfüoğlu (2025a); Bolokhov and Skvortsova (2025c); Dubinsky (2025a); Malik (2025b); Bolokhov and Skvortsova (2025b); Dubinsky (2025b); Lütfüoğlu (2025b); Malik (2025c); Lütfüoğlu et al. (2025b). In the present work, we test this correspondence explicitly for higher-dimensional regular black holes in quasi-topological gravity by comparing the behavior of grey-body factors with the analytic approximation based on the quasinormal frequencies obtained for the same backgrounds.
@@
This paper is designed as follows. In Sec. II, we introduce the regular black hole configurations in quasi-topological gravity and the master equations governing electromagnetic perturbations. In Sec. III, we briefly review the WKB approach used to compute the grey-body factors. In Sec. IV, we use the numerically obtained grey-body factors to evaluate the energy-emission rates and discuss the evaporation dynamics. Finally, in Sec. V, we summarize our results and discuss their physical implications.
II Maxwell field in the background of the quasi-topological black holes
Label $f(r)$ $\mu$ parametter in horizon units Constraint in $\alpha$ Hawking temperature $(4\pi T_{H})$
$(a)$ $1-\frac{\mu r^{2}}{r^{D-1}+\alpha\mu}$ $\frac{r_{0}^{D-1}}{r_{0}^{2}-\alpha}$ $0\leq\frac{\alpha}{r_{0}^{2}}\leq\frac{D-3}{D-1}$ $\frac{1}{r_{0}^{3}}\left[(D-3)r_{0}^{2}-(D-1)\alpha\right]$
$(b)$ $1-\frac{\mu r^{2}}{\sqrt{r^{2(D-1)}+\alpha^{2}\mu^{2}}}$ $\frac{r_{0}^{D-1}}{\sqrt{r_{0}^{4}-\alpha^{2}}}$ $0\leq\frac{\alpha}{r_{0}^{2}}\leq\sqrt{\frac{D-3}{D-1}}$ $\frac{1}{r_{0}^{5}}\left[(D-3)r_{0}^{4}-(D-1)\alpha^{2}\right]$
$(c)$ $1-\frac{r^{2}}{\alpha}\left(1-e^{-\alpha\mu/r^{D-1}}\right)$ $\frac{r_{0}^{D-1}}{\alpha}\ln{\frac{r_{0}^{2}}{r_{0}^{2}-\alpha}}$ $0\leq\frac{\alpha}{r_{0}^{2}}\leq C(D)<1$ $\frac{1}{r_{0}\alpha}\left[(D-1)(r_{0}^{2}-\alpha)\ln{\left|\frac{r_{0}^{2}}{r_{0}^{2}-\alpha}\right|}-2\alpha\right]$
$(d)$ $1-\frac{2\mu r^{2}}{r^{D-1}+\sqrt{r^{2(D-1)}+4\alpha^{2}\mu^{2}}}$ $\frac{r_{0}^{D+1}}{r_{0}^{4}-\alpha^{2}}$ $0\leq\frac{\alpha}{r_{0}^{2}}\leq\sqrt{\frac{D-3}{D+1}}$ $\frac{(D-3)r_{0}^{4}-(D+1)\alpha^{2}}{r_{0}(r_{0}^{4}+\alpha^{2})}$
$(e)$ $1-\frac{2\mu r^{2}}{r^{D-1}+2\alpha\mu+\sqrt{r^{2(D-1)}+4\mu\alpha r^{D-1}}}$ $\frac{r_{0}^{D+1}}{(r_{0}^{2}-\alpha)^{2}}$ $0\leq\frac{\alpha}{r_{0}^{2}}\leq\frac{D-3}{D+1}$ $\frac{(D-3)r_{0}^{2}-(D+1)\alpha}{r_{0}(r_{0}^{2}+\alpha)}$
$(f)$ $1-\frac{2\mu r^{2}}{\mu\alpha+\sqrt{4r^{2(D-1)}+\mu^{2}\alpha^{2}}}$ $\frac{r_{0}^{D-2}}{\sqrt{r_{0}^{2}-\alpha}}$ $0\leq\frac{\alpha}{r_{0}^{2}}\leq\frac{D-3}{D-2}$ $\frac{2}{r_{0}}\frac{(D-3)r_{0}^{2}-(D-2)\alpha}{2r_{0}^{2}-\alpha}$
Table 1: Summary of the considered regular black hole models. Configurations (a)(a)$(e)$ were proposed in Bueno et al. (2025a), while configuration $(f)$ was introduced in Arbelaez (2025). The constraint on $\alpha$ for configuration $(c)$, denoted $C(D)$, does not admit a closed-form expression. Its value increases monotonically with $D$, $C(5)\approx 0.715$, $C(6)\approx 0.802$, and approaches unity asymptotically, $\lim_{D\to\infty}C(D)=1$.
@@ II Maxwell field in the background of the quasi-topological black holes
Following Ref. Bueno et al. (2020), we consider a higher-curvature gravitational theory whose dynamics are governed by the action
$I_{QT}=\frac{1}{16\pi G}\int{d^{D}x\sqrt{|g|}\left[R+\sum^{n_{max}}_{n=2}{\alpha_{n}\mathcal{Z}_{n}}\right]},$ (1)
@@
The static, spherically symmetric solutions of this theory can be written in the form Bueno et al. (2025a):
$ds^{2}=-N(r)^{2}f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2}_{D-2},$ (2)
@@
$h(\psi)\equiv\psi+\sum_{n=2}^{n_{max}}{\alpha_{n}\psi^{n}},\quad h(\psi)\equiv\frac{\mu}{r^{D-1}},$ (4)
@@
$\mu\equiv\frac{16\pi G_{D}M}{(D-2)\Omega_{D-2}},\quad\Omega_{D-2}=\frac{(2\pi)^{\frac{D-1}{2}}}{\Gamma(\frac{D-1}{2})}.$ (5)
For appropriate choices of the couplings $\alpha_{n}$, this framework admits regular black hole solutions in which curvature invariants remain finite everywhere. Table 1 summarizes the different regular black hole geometries arising from infinite–curvature corrections that will be considered in this work. These spacetimes serve as background geometries for the computation of grey-body factors and the associated Hawking energy emission spectra.
@@
$\frac{1}{\sqrt{-g}}\partial_{\mu}\Big(\sqrt{-g}\,F_{\rho\sigma}g^{\rho\nu}g^{\sigma\mu}\Big)$ $=0,$ (6)
@@
Following Ref. Crispino et al. (2001), we adopt Feynman’s gauge, which allows a separation of variables. After decomposing the vector potential into scalar-type and vector-type components on the (D2)(D-2)-sphere, the perturbation equations reduce to a set of one–dimensional wave equations of Schrödinger type,
$\frac{d^{2}\Psi(r_{*})}{dr_{*}^{2}}+\left[\omega^{2}-V(r_{*})\right]\Psi(r_{*})=0,$ (7)
@@
The effective potentials governing the scalar-type and vector-type electromagnetic perturbations are given, respectively, by $V_{S}(r)$ and $V_{V}(r)$ Lopez-Ortega (2006):
$V_{S}(r)=f(r)\left(\frac{\ell(\ell+D-3)}{r^{2}}+\right.\\ \frac{(D-2)(D-4)}{4r^{2}}f(r)\left.-\frac{D-4}{2r}\frac{df}{dr}\right),$ (8a)
$V_{V}(r)=f(r)\left(\frac{(\ell+1)(\ell+D-4)}{r^{2}}+\right.\\ \frac{(D-4)(D-6)}{4r^{2}}f(r)\left.+\frac{D-4}{2r}\frac{df}{dr}\right),$ (8b)
@@
III WKB approach for the calculation of grey-body factors
We consider the scattering problem associated with Eq. (7), subject to the boundary conditions
$\Psi$ $=e^{-i\omega r_{*}}+R\,e^{i\omega r_{*}},$ $r_{*}\to+\infty,$ (9a)
$\Psi$ $=T\,e^{-i\omega r_{*}},$ $r_{*}\to-\infty.$ (9b)

### v2

@@ 1 Introduction
Black holes are not perfect black bodies. Quantum mechanically, they emit radiation with a thermal spectrum modulated by frequency-dependent transmission coefficients, known as grey-body factors, which encode the interaction of the emitted quanta with the curved spacetime geometry outside the event horizon. These factors determine both the energy emission rates and the detailed spectral properties of Hawking radiation [45], thereby providing a direct link between semiclassical quantum effects and the classical scattering problem for perturbations in black hole backgrounds. As such, grey-body factors play a central role in understanding black hole evaporation, stability, and potential observational signatures of modifications to General Relativity [102, 103, 50].
@@
In recent years, regular black holes – spacetimes free of curvature singularities – have attracted considerable attention as viable effective descriptions of quantum-corrected gravitational collapse. In many constructions, regularity is achieved either through effective matter sources or via higher-curvature corrections, while preserving the existence of an event horizon. From a phenomenological perspective, such models provide a setting in which quantum-gravity-inspired modifications can be explored without encountering pathological divergences at the classical level. However, regularity alone does not guarantee that the dynamical, thermodynamical, or radiative properties of these black holes closely resemble those of their singular counterparts. This makes a detailed analysis of wave propagation, scattering, and Hawking emission indispensable. Consequently, a substantial body of literature has been devoted to the study of perturbations, quasinormal spectra, grey-body factors, and Hawking radiation of various regular black hole models [108, 78, 36, 96, 76, 55, 70, 100, 38, 17, 3, 53, 74, 49, 6, 46, 14, 2, 42, 115, 105, 40, 9, 111, 18, 110, 64, 16, 12, 11, 84, 25, 26, 98].
@@
Higher-curvature extensions of General Relativity are particularly well suited for such investigations. Among them, quasi-topological gravity occupies a distinguished position, as it allows for nontrivial higher-order curvature corrections while retaining second-order field equations for static, spherically symmetric spacetimes in arbitrary dimensions [19, 113, 24, 28]. When combined with regular black hole metrics, quasi-topological gravity provides a consistent and tractable framework in which higher-curvature effects and regular cores jointly influence black hole radiation and dynamical properties [22, 62, 61, 20, 23, 39]. Notice, that black holes in theories with usual Gauss-Bonnet or Lovelock corrections are usually singular and their quasinormal spectrum has been extensively studied (see, for instance [1, 99, 32, 117, 114, 31, 56, 57, 51, 58, 41, 118, 29, 66, 27, 7, 65, 8, 106, 71] and references therein).
@@
In this work, we investigate grey-body factors and energy emission rates for higher-dimensional regular black holes arising in quasi-topological gravity. By solving the wave equations for the Maxwell field propagating in these backgrounds, we analyze how spacetime dimensionality and regularization parameter affect transmission probabilities and Hawking spectra. Special attention is devoted to deviations from purely thermal emission and to the role of the effective potential barrier surrounding the black hole. While Hawking radiation for two specific regular black hole solutions in quasi-topological gravity was previously considered in Ref. [75] for particular values of the coupling parameter, a comprehensive analysis covering a broader class of regular solutions and the full range of parameters has not yet been performed. The present work aims to fill this gap through a systematic study of grey-body factors and energy emission rates for regular black holes in quasi-topological gravity.
@@
In addition, we explore the correspondence between quasinormal modes and grey-body factors, treating it as a complementary aspect rather than a primary assumption. Quasinormal modes govern the characteristic damped oscillations of black holes and are closely related to the poles of scattering amplitudes in the complex frequency plane. It has been suggested in various contexts that features of the grey-body spectrum, such as characteristic frequency scales or transmission resonances, may be linked to the quasinormal mode spectrum [90, 44, 87, 10, 35, 89, 11, 34, 80, 91, 85]. In the present work, we test this correspondence explicitly for higher-dimensional regular black holes in quasi-topological gravity by comparing the behavior of grey-body factors with the analytic approximation based on the quasinormal frequencies obtained for the same backgrounds.
@@
This paper is designed as follows. In Sec. 2, we introduce the regular black hole configurations in quasi-topological gravity and the master equations governing electromagnetic perturbations. In Sec. 3, we briefly review the WKB approach used to compute the grey-body factors. In Sec. 4, we use the numerically obtained grey-body factors to evaluate the energy-emission rates and discuss the evaporation dynamics. Finally, in Sec. 5, we summarize our results and discuss their physical implications.
2 Maxwell field in the background of the quasi-topological black holes
@@
Following Ref. [21], we consider a higher-curvature gravitational theory whose dynamics are governed by the action
$I_{QT}=\frac{1}{16\pi G}\int{d^{D}x\sqrt{|g|}\left[R+\sum^{n_{max}}_{n=2}{\alpha_{n}\mathcal{Z}_{n}}\right]},$ (2.1)
@@
The static, spherically symmetric solutions of this theory can be written in the form [22]:
$ds^{2}=-N(r)^{2}f(r)dt^{2}+\frac{dr^{2}}{f(r)}+r^{2}d\Omega^{2}_{D-2},$ (2.2)
@@
$h(\psi)\equiv\psi+\sum_{n=2}^{n_{max}}{\alpha_{n}\psi^{n}},\quad h(\psi)\equiv\frac{\mu}{r^{D-1}},$ (2.4)
@@
$\mu\equiv\frac{16\pi G_{D}M}{(D-2)\Omega_{D-2}},\quad\Omega_{D-2}=\frac{(2\pi)^{\frac{D-1}{2}}}{\Gamma(\frac{D-1}{2})}.$ (2.5)
For appropriate choices of the couplings $\alpha_{n}$, this framework admits regular black hole solutions in which curvature invariants remain finite everywhere.
Label $f(r)$ $\mu$ parameter in horizon units Constraint in $\alpha$
$(a)$ $1-\frac{\mu r^{2}}{r^{D-1}+\alpha\mu}$ $\frac{r_{0}^{D-1}}{r_{0}^{2}-\alpha}$ $0\leq\frac{\alpha}{r_{0}^{2}}\leq\frac{D-3}{D-1}$
$(b)$ $1-\frac{\mu r^{2}}{\sqrt{r^{2(D-1)}+\alpha^{2}\mu^{2}}}$ $\frac{r_{0}^{D-1}}{\sqrt{r_{0}^{4}-\alpha^{2}}}$ $0\leq\frac{\alpha}{r_{0}^{2}}\leq\sqrt{\frac{D-3}{D-1}}$
$(c)$ $1-\frac{r^{2}}{\alpha}\left(1-e^{-\alpha\mu/r^{D-1}}\right)$ $\frac{r_{0}^{D-1}}{\alpha}\ln\!\left(\frac{r_{0}^{2}}{r_{0}^{2}-\alpha}\right)$ $0\leq\frac{\alpha}{r_{0}^{2}}\leq C(D)<1$
$(d)$ $1-\frac{2\mu r^{2}}{r^{D-1}+\sqrt{r^{2(D-1)}+4\alpha^{2}\mu^{2}}}$ $\frac{r_{0}^{D+1}}{r_{0}^{4}-\alpha^{2}}$ $0\leq\frac{\alpha}{r_{0}^{2}}\leq\sqrt{\frac{D-3}{D+1}}$
$(e)$ $1-\frac{2\mu r^{2}}{r^{D-1}+2\alpha\mu+\sqrt{r^{2(D-1)}+4\mu\alpha r^{D-1}}}$ $\frac{r_{0}^{D+1}}{(r_{0}^{2}-\alpha)^{2}}$ $0\leq\frac{\alpha}{r_{0}^{2}}\leq\frac{D-3}{D+1}$
$(f)$ $1-\frac{2\mu r^{2}}{\mu\alpha+\sqrt{4r^{2(D-1)}+\mu^{2}\alpha^{2}}}$ $\frac{r_{0}^{D-2}}{\sqrt{r_{0}^{2}-\alpha}}$ $0\leq\frac{\alpha}{r_{0}^{2}}\leq\frac{D-3}{D-2}$
Table 1: Summary of the considered regular black hole models. Configurations (a)(a)$(e)$ were proposed in [22], while configuration $(f)$ was introduced in [5]. The constraint on $\alpha$ for configuration $(c)$, denoted $C(D)$, does not admit a closed-form expression. Its value increases monotonically with $D$, $C(5)\approx 0.715$, $C(6)\approx 0.802$, and approaches unity asymptotically, $\lim_{D\to\infty}C(D)=1$.
Tables 1 and 2 summarize the different regular black hole geometries arising from infinite–curvature corrections that will be considered in this work. These spacetimes serve as background geometries for the computation of grey-body factors and the associated Hawking energy emission spectra.
Label $h(\psi)$ $4\pi T_{H}$
$(a)$ $\frac{\psi}{1-\alpha\psi}$ $\frac{1}{r_{0}^{3}}\left[(D-3)r_{0}^{2}-(D-1)\alpha\right]$
$(b)$ $\frac{\psi}{\sqrt{1-\alpha^{2}\psi^{2}}}$ $\frac{1}{r_{0}^{5}}\left[(D-3)r_{0}^{4}-(D-1)\alpha^{2}\right]$
$(c)$ $-\frac{\log{\left(1-\alpha\psi\right)}}{\alpha}$ $\frac{1}{r_{0}\alpha}\left[(D-1)(r_{0}^{2}-\alpha)\ln\!\left|\frac{r_{0}^{2}}{r_{0}^{2}-\alpha}\right|-2\alpha\right]$
$(d)$ $\frac{\psi}{1-\alpha^{2}\psi^{2}}$ $\frac{(D-3)r_{0}^{4}-(D+1)\alpha^{2}}{r_{0}(r_{0}^{4}+\alpha^{2})}$
$(e)$ $\frac{\psi}{\left(1-\alpha\psi\right)^{2}}$ $\frac{(D-3)r_{0}^{2}-(D+1)\alpha}{r_{0}(r_{0}^{2}+\alpha)}$
$(f)$ $\frac{\psi}{\sqrt{1-\alpha\psi}}$ $\frac{2}{r_{0}}\frac{(D-3)r_{0}^{2}-(D-2)\alpha}{2r_{0}^{2}-\alpha}$
Table 2: Characteristic functions $h(\psi)$ and Hawking temperatures for the regular black hole models considered in this work.
@@
$\frac{1}{\sqrt{-g}}\partial_{\mu}\Big(\sqrt{-g}\,F_{\rho\sigma}g^{\rho\nu}g^{\sigma\mu}\Big)$ $=0,$ (2.6)
@@
Following Ref. [30], we adopt Feynman’s gauge, which allows a separation of variables. After decomposing the vector potential into scalar-type and vector-type components on the (D2)(D-2)-sphere, the perturbation equations reduce to a set of one–dimensional wave equations of Schrödinger type,
$\frac{d^{2}\Psi(r_{*})}{dr_{*}^{2}}+\left[\omega^{2}-V(r_{*})\right]\Psi(r_{*})=0,$ (2.7)
@@
The effective potentials governing the scalar-type and vector-type electromagnetic perturbations are given, respectively, by $V_{S}(r)$ and $V_{V}(r)$ [79]:
$V_{S}(r)$ $=f(r)\left(\frac{\ell(\ell+D-3)}{r^{2}}+\frac{(D-2)(D-4)}{4r^{2}}f(r)-\frac{D-4}{2r}\frac{df}{dr}\right),$ (2.8a)
$V_{V}(r)$ $=f(r)\left(\frac{(\ell+1)(\ell+D-4)}{r^{2}}+\frac{(D-4)(D-6)}{4r^{2}}f(r)+\frac{D-4}{2r}\frac{df}{dr}\right),$ (2.8b)
@@
3 WKB approach for the calculation of grey-body factors
We consider the scattering problem associated with Eq. (2.7), subject to the boundary conditions
$\Psi$ $=e^{-i\omega r_{*}}+R\,e^{i\omega r_{*}},$ $r_{*}\to+\infty,$ (3.1a)
$\Psi$ $=T\,e^{-i\omega r_{*}},$ $r_{*}\to-\infty.$ (3.1b)